# Enhancing the Sensitivity to Seesaw Predictions in Gauged $B-L$ Scenarios

### v1

@@ I Introduction
The observation of neutrino flavor oscillations implies the existence of neutrino masses. While a large number of possible scenarios can account for neutrino masses, all of them require physics beyond the Standard Model (SM). A particularly well-studied possibility is the so-called “seesaw” scenario in which new right-handed sterile neutrinos mass mix with the left-handed neutrinos in such a way as to explain the smallness of the neutrino masses Minkowski:1977sc ; Gell-Mann:1979vob ; Mohapatra:1979ia . The existence of such states may be further motivated by new gauge symmetries. In particular, the combination of baryon and lepton number, $B-L$, is non-anomalous in the SM, provided there exist three additional right-handed neutrinos Mohapatra:1980qe , making it a favorable extension of the SM. Thus, gauging the combination of $B-L$ can provide yet additional motivation, beyond neutrino masses, for the necessity of right-handed sterile neutrinos.
@@ II Production channels of $Z^{\prime}$
In proton beam-dump experiments the new $B-L$ gauge boson can be produced either directly via proton bremsstrahlung or indirectly via secondary SM particles. The secondary particles we are interested in are photons and positrons. In the context of $B-L$, each photon can be substituted by a $Z^{\prime}$ when kinematics allows it. In particular, we consider photons from neutral meson decays and electron bremsstrahlung. On the other hand, positrons can lead to resonant production of $Z^{\prime}$ through on-shell annihilation with electrons. In general, we find that the contribution from meson decays is almost negligible, but it provides a useful comparison with the results available in the literature. Here we do not consider $Z^{\prime}$ production through Compton scattering Celentano:2020vtu , since it does not change our conclusions in the relevant region of the parameter space. We also neglect charged meson decays, but in § V, we compare our final results with those obtained in ref.Berryman:2019dme , where such a channel is studied in detail.
@@ II.1 Proton Bremsstrahlung
Figure 1: Proton bremsstrahlung via Pomeron exchange, producing the on-shell $U(1)_{B-L}$ gauge boson that promptly decays to $NN$ pairs.
First, we consider the $Z^{\prime}$ bremsstrahlung of the primary beam protons, Fig. 1. This process is modeled by the quasi-real approximation in non-single diffractive Pomeron exchange, calculated in ref. Foroughi-Abari:2021zbm . We briefly summarize the calculation as follows. The differential scattering cross-section is given by
$\dfrac{\partial^{2}\sigma(pp\to ppZ^{\prime})}{\partial p_{T}^{2}\partial z}=w^{D}(z,p_{T}^{2})\times\sigma_{pp}^{\text{NSD}}(s^{\prime})$ (2)
where $w^{D}(z,p_{T}^{2})$ is a splitting function over the longitudinal momentum fraction $z=p_{Z^{\prime}}/p$ of the outgoing $Z^{\prime}$ momentum with respect to the initial proton momentum $p$, $p_{T}$ is the transverse momentum of the outgoing $Z^{\prime}$, and $s^{\prime}=2m_{p}(p(1-z)+m_{p})$ where $m_{p}$ is the proton mass. In the integration of Eq. 2, certain cuts have to be made to ensure the applicability of the quasi-real approximation; following ref. Foroughi-Abari:2021zbm , we define
@@
Assuming that most of the protons get absorbed in the proton target, we can estimate the differential rate of $Z^{\prime}$ particles produced by proton bremsstrahlung as the number of POT times the ratio of the bremsstrahlung and total proton cross sections Berryman:2019dme ; Dev:2023zts ; [^1] During the completion of this work, another analysis of proton bremsstrahlung in the quasi-real approximation was studied in ref. Foroughi-Abari:2024xlj which included an improved treatment of the form factors and the treatment of $\phi$-meson-nucleon interactions, among others. We do not expect our results will change qualitatively after using the updated cross sections in ref. Foroughi-Abari:2024xlj , except in a narrow range within $m_{Z^{\prime}}\simeq 1$ GeV, where the $\phi$ mass resonance contributes an enhancement to the $Z^{\prime}$ production cross section. This possibility will be considered in a future revision of the present work.
$\dfrac{\partial^{2}N_{Z^{\prime}}}{\partial p_{T}^{2}\partial z}=N_{\text{POT}}\times\dfrac{1}{\sigma_{\text{tot}}(s)}\times\dfrac{\partial^{2}\sigma(pp\to ppZ^{\prime})}{\partial p_{T}^{2}\partial z}\times\Theta_{cuts}$ (4)
@@ II.4 Resonant Production
We show the rates of $Z^{\prime}$ that are produced and have 3-momentum that points within the solid angle of the DUNE near detector in Fig. 4 for resonant production, electron/positron bremsstrahlung, neutral meson decay ($\pi^{0},\eta$), and proton bremsstrahlung as a function of the $Z^{\prime}$ mass. We find that the electron/positron channels contribute mainly to the low mass limit, while proton bremsstrahlung dominates the production for 100 MeV - 1 GeV masses.
@@ IV Main results
Figure 8: Number of particles from all visible decays in DUNE as a function of $m_{N}$ and $\left|U_{\mu}\right|^{2}$ (left column) and $\left|U_{\tau}\right|^{2}$ (right column). Top panels refer to $\frac{m_{Z^{\prime}}}{m_{N}}=2.1$ while bottom ones refer to $\frac{m_{Z^{\prime}}}{m_{N}}=5$. We fix the coupling of the $Z^{\prime}$ to $g=10^{-4}$. The solid, dashed, and dotted lines correspond to the iso-event contours for 3, 10, and 100 events, respectively. We also show the forecasted DUNE-ND sensitivity to HNLs produced only from the $|U_{\mu}|$ mixing-angle driven channels (teal dotted line) Krasnov:2019kdc ; Ballett:2019bgd .
@@ Acknowledgements
AT acknowledges support in part by the US Department of Energy (DOE) grant #DE-SC0010143. IMS is supported by the US Department of Energy under the award numbers DE-SC0020250 and DE-SC0020262. The work of F.C. was partially supported by the research grant number 2022E2J4RK “PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR) and by the European Union - Next Generation EU. The work of BD is supported in part by the U.S. Department of Energy Grant DE-SC0010813. The work of GG, WJ and JY is supported by the U.S. Department of Energy under Grant No. DE-SC0011686.

### v2

@@ I Introduction
The observation of neutrino flavor oscillations implies the existence of neutrino masses. While a large number of possible scenarios can account for neutrino masses, all of them require physics beyond the Standard Model (SM). A particularly well-studied possibility is the so-called “seesaw” scenario in which new right-handed sterile neutrinos mass mix with the left-handed neutrinos in such a way as to explain the smallness of the neutrino masses Minkowski:1977sc ; Gell-Mann:1979vob ; Mohapatra:1979ia . The existence of such states may be further motivated by new gauge symmetries. In particular, the combination of baryon and lepton number, $B-L$, is non-anomalous in the SM, provided there exist three additional right-handed neutrinos PhysRevD.20.776 ; Mohapatra:1980qe , making it a favorable extension of the SM. Thus, gauging the combination of $B-L$ can provide yet additional motivation, beyond neutrino masses, for the necessity of right-handed sterile neutrinos.
@@ II Production channels of $Z^{\prime}$
In proton beam-dump experiments the new $B-L$ gauge boson can be produced either directly via proton bremsstrahlung or indirectly via secondary SM particles. The secondary particles we are interested in are photons and positrons. In the context of $B-L$, each photon can be substituted by a $Z^{\prime}$ when kinematics allows it. In particular, we consider photons from neutral meson decays and electron bremsstrahlung. On the other hand, positrons can lead to resonant production of $Z^{\prime}$ through on-shell annihilation with electrons. Here we do not consider $Z^{\prime}$ production through Compton scattering Celentano:2020vtu , since it has a smaller cross section relative to the other processes and does not change our conclusions in the relevant region of the parameter space. We also neglect charged meson decays, but in § V, we compare our final results with those obtained in ref.Berryman:2019dme , where such a channel is studied in detail.
@@ II.1 Proton Bremsstrahlung
Figure 1: Proton bremsstrahlung via non-single diffractive scatttering, producing the on-shell $U(1)_{B-L}$ gauge boson that promptly decays to $NN$ pairs.
First, we consider the $Z^{\prime}$ bremsstrahlung of the primary beam protons, Fig. 1. The diffractive scattering process is modeled in the quasi-real approximation by considering producing the vector $Z^{\prime}$ as initial state radiation (see ref. Foroughi-Abari:2021zbm for details). We briefly summarize the calculation as follows. The differential scattering cross-section is given by
$\dfrac{\partial^{2}\sigma(pp\to XZ^{\prime})}{\partial p_{T}^{2}\partial z}=w^{D}(z,p_{T}^{2})\times\sigma_{pp}^{\text{NSD}}(s^{\prime})$ (2)
where $w^{D}(z,p_{T}^{2})$ is a splitting function over the longitudinal momentum fraction $z=p_{Z^{\prime}}/p$ of the outgoing $Z^{\prime}$ momentum with respect to the initial proton momentum $p$, $p_{T}$ is the transverse momentum of the outgoing $Z^{\prime}$, and $s^{\prime}=2m_{p}(p(1-z)+m_{p})$ where $m_{p}$ is the proton mass. In the construction of the splitting function $w^{D}(z,p_{T}^{2})$ in ref. Foroughi-Abari:2021zbm , there are hadronic form factors introduced to capture timelike momentum transfer as well as the off-shell momentum in Fig. 1. The parametrization of the form factors can have a significant impact on the magnitude of the cross section (up to an order of magnitude), especially in the parameter $\Lambda_{p}\sim\mathcal{O}(m_{p})$ which acts as a cutoff scale for the off-shell behavior. In this work we take the central value in the range considered by ref. Foroughi-Abari:2021zbm for the cutoff scale, $\Lambda_{p}=1.5$ GeV.
In the integration of Eq. 2, certain cuts have to be made to ensure the applicability of the quasi-real approximation; following ref. Foroughi-Abari:2021zbm , we define
@@
Assuming that most of the protons get absorbed in the proton target, we can estimate the differential rate of $Z^{\prime}$ particles produced by proton bremsstrahlung as the number of POT times the ratio of the bremsstrahlung and total proton cross sections Berryman:2019dme ; Dev:2023zts ; [^1] During the completion of this work, other analyses have proposed updated treatments of the form factors used proton bremsstrahlung, see for example ref. Foroughi-Abari:2024xlj and refs. Gorbunov:2024vrc ; Gorbunov:2023jnx . The potential impact of these new calculations is left to a future work.
$\dfrac{\partial^{2}N_{Z^{\prime}}}{\partial p_{T}^{2}\partial z}=N_{\text{POT}}\times\dfrac{1}{\sigma_{\text{tot}}(s)}\times\dfrac{\partial^{2}\sigma(pp\to XZ^{\prime})}{\partial p_{T}^{2}\partial z}\times\Theta_{cuts}$ (4)
@@ II.4 Resonant Production
We show the rates of $Z^{\prime}$ that are produced and have 3-momentum that points within the solid angle of the DUNE near detector in Fig. 4 for resonant production, electron/positron bremsstrahlung, neutral meson decay ($\pi^{0},\eta$), and proton bremsstrahlung as a function of the $Z^{\prime}$ mass. We find that the electron/positron channels contribute mainly to the low mass limit, while proton bremsstrahlung dominates the production for 250 MeV - 1 GeV masses. The extrapolated cross section for proton bremsstrahlung does grow at lower masses, shown by the dashed line in Fig. 4, but we only consider this process above $m_{Z^{\prime}}>250$ MeV, below which the appropriate behavior becomes theoretically uncertain.
@@ IV Main results
Figure 8: Number of particles from all visible decays in DUNE as a function of $m_{N}$ and $\left|U_{\mu}\right|^{2}$ (left column) and $\left|U_{\tau}\right|^{2}$ (right column). Top panels refer to $\frac{m_{Z^{\prime}}}{m_{N}}=2.1$ while bottom ones refer to $\frac{m_{Z^{\prime}}}{m_{N}}=5$. We fix the coupling of the $Z^{\prime}$ to $g=10^{-4}$. The solid, dashed, and dotted lines correspond to the iso-event contours for 3, 10, and 100 events, respectively, separated by production channel (neutral meson decays, electron/positron bremsstrahlung and resonant production, and proton bremsstrahlung). We also show the forecasted DUNE-ND sensitivity to HNLs produced only from the $|U_{\mu}|$ mixing-angle driven channels (teal dotted line) Krasnov:2019kdc ; Ballett:2019bgd .
@@ Acknowledgements
We thank Kevin J. Kelly for the useful feedback and discussions. We also thank Ekaterina Kriukova and Dmitry Gorbunov for their helpful comments. AT acknowledges support in part by the US Department of Energy (DOE) grant #DE-SC0010143. IMS is supported by the US Department of Energy under the award numbers DE-SC0020250 and DE-SC0020262. The work of F.C. was partially supported by the research grant number 2022E2J4RK “PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR) and by the European Union - Next Generation EU. The work of BD is supported in part by the U.S. Department of Energy Grant DE-SC0010813. The work of GG, WJ and JY is supported by the U.S. Department of Energy under Grant No. DE-SC0011686.